### v1

# Large Language Model Performance Benchmarking on Mobile Platforms: A Thorough Evaluation

Abstract
As large language models (LLMs) increasingly integrate into every aspect of our work and daily lives, there are growing concerns about user privacy, which push the trend toward local deployment of these models. There are a number of lightweight LLMs (e.g., Gemini Nano, LLAMA2 7B) that can run locally on smartphones, providing users with greater control over their personal data. As a rapidly emerging application, we are concerned about their performance on commercial-off-the-shelf mobile devices. To fully understand the current landscape of LLM deployment on mobile platforms, we conduct a comprehensive measurement study on mobile devices. We evaluate both metrics that affect user experience, including token throughput, latency, and battery consumption, as well as factors critical to developers, such as resource utilization, DVFS strategies, and inference engines. In addition, we provide a detailed analysis of how these hardware capabilities and system dynamics affect on-device LLM performance, which may help developers identify and address bottlenecks for mobile LLM applications. We also provide comprehensive comparisons across the mobile system-on-chips (SoCs) from major vendors, highlighting their performance differences in handling LLM workloads. We hope that this study can provide insights for both the development of on-device LLMs and the design for future mobile system architecture.
1 Introduction
Over the past two years, Large Language Models (LLMs) have demonstrated exceptional performance across a wide range of domains. Their advanced capabilities in understanding human language and generating text have made them indispensable in various applications, including dialogue systems, machine translation, and text generation. Several companies have begun integrating LLMs as system services (e.g., Apple Intelligence [21], Copilot+ PC [32]), where these models assist users in organizing daily routines, summarizing emails and messages, and even generating automated replies. However, while LLMs serve as powerful and appealing personal assistants, these models continuously monitor and analyze users’ activities, which leads to serious privacy concerns [19].
To address the privacy concerns, local deployment of LLMs is becoming attractive, as it allows these models to perform inference directly on the device and no personal data are sent to cloud servers. In this way, users have greater control of their personal data. While LLMs, such as ChatGPT [2] and Claude [4], are typically deployed on the cloud, local deployment are becoming feasible for two reasons. On one hand, a series of lightweight LLMs (e.g., Llama2-7B [41] and Mistral-7B [22]) have been introduced. These models, with billions of parameters, are significantly smaller in size compared to cloud-based models, which typically have hundreds of billions of parameters. Despite their reduced size, these lightweight LLMs maintain a high level of expressiveness in both common-sense reasoning and domain-specific knowledge, and thus can fully satisfy the basic needs of mobile users. On the other hand, state-of-the-art mobile SoCs demonstrate significant improvement in computing capability and memory capacity. More importantly, a variety of specialized AI accelerators are integrated into the mobile processors, such as Graphics Processing Units (GPUs) and Neural Processing Units (NPUs). These hardware improvements make it possible to implement complex AI functions on mobile devices.
As a rapidly emerging application, LLMs differ significantly from traditional mobile applications in terms of computational complexity and memory demands. Therefore, we are particularly concerned with their performance on commercial off-the-shelf (COTS) mobile devices, as well as the performance fluctuations caused by mobile operating systems and hardware architecture compatibility. Several studies have taken the first steps to explore the current state of mobile LLM[25][27][34]. They have conducted preliminary assessments of model deployment and inference performance through the development of specialized benchmark suites. However, they mainly focus on performance metrics such as token throughput and latency, without looking into the underlying hardware dynamics and system configurations, such as CPU/GPU utilization, number of threads, operating frequencies, etc. Besides, there is a lack of analysis about how these hardware and system-level factors affect on-device LLM inference. This is essential for developers to identify and address bottleneck for mobile LLM applications.
In this paper, we aim to reveal the current state and potential of mobile hardware for supporting on-device LLMs by deploying LLMs across a diverse range of mobile devices. We select multiple high-tier SoCs from different vendors that represent currently available resources and development trends in mobile hardware. We analyze the differences in SoCs from major vendors and ultimately select SoCs represented by Qualcomm, HiSilicon, and MediaTek to perform local LLM inference. We deploy a B model on mobile devices with llama.cpp [14] and MLC LLM [38], which are the two popular mobile LLM inference engines. During inference, we collect comprehensive metrics with specific profilers including Snapdragon Profiler [35] and Arm Streamline [5] to make sure that all the data accurately reflect the hardware performance. We investigate how the hardware specifications, including big.LITTLE cores[6], cache size, and memory bandwidth, affect the inference performance, which is essential to help the development of future mobile SoCs to accelerate LLMs.
From our study, we have the following key observations:
1.
While LLMs are typically considered to be resource-intensive, the current implementation does not fully explore the potential of mobile processors. It only utilizes of the arithmetic units on mobile GPUs.
2.
Processors with superior hardware specifications may not necessarily exhibit better performance (e.g., Mali-G720 Immortalis MP12 vs Adreno 750). In addition, the Adreno GPUs consistently outperform the Mali GPUs in overall performance.
3.
Given the "big.LITTLE" architecture of mobile CPUs, in most cases, setting the number of threads equal to the number of big cores (prime and performance cores) can achieve the optimal performance. Adding more cores (efficiency cores) may degrade the performance.
4.
Specialized machine instructions, such as smmla and sdot in Arm architecture, significantly enhance LLM performance. By carefully rearranging the weight matrix, these instructions can yield up to a speed improvement.
5.
Specialized AI accelerators, such as the Hexagon NPU, demonstrate a remarkable prefill speed, representing a -fold increase compared to CPU and GPU. However, their performance in decoding does not exhibit a similar level of enhancement.
We summarize our contributions as follows:
1.
We present a comprehensive measurement study of LLM inference on mobile platforms. We present not only metrics relevant to user experience, but also hardware and system characteristics that are critical to developers.
2.
We provide a thorough analysis of how hardware attributes and system dynamics impact on-device LLM performance and highlights potential bottlenecks in mobile LLM applications.
3.
Based on our measurement results, we propose several potential directions to accelerate on-device LLM inference.
The rest of the paper is organized as follows. In Section 2, we present preliminaries and related works on this topic. In Section 3, we describe our measurement setup and methodology. In Section 4, we present performance metrics that are of primary concern to users, while Section 5 discusses findings related to the underlying hardware characteristics and inference framework, which may interest developers. We discuss potential improvements in Section 6 and explain the remaining issues in Section 7. Finally, we conclude the paper in Section 8.
2 Related Works
2.1 Light-Weight LLMs
The number of LLMs has increased significantly in recent years. Numerous studies have explored the potential performance limits of these models by leveraging scaling laws [23] to increase their size. However, given the severe constraints of mobile platforms, it is crucial to compress these models to ensure that they are lightweight enough for deployment on edge devices. Current methods for compressing large models usually fall into the categories of pruning [29][42], distillation [40], quantization [13][28], and low-order decomposition [45]. Quantization is a widely used resource-efficient technique for compressing LLMs, reducing their memory and bandwidth requirements. The nature of quantization is a mapping from floating-point numbers to integers, especially the 4-bit integer quantization for mobile LLMs, which is recognized as the optimal compromise between model size and accuracy [27]. For example, based on the second-order information for weight updating, GPTQ [13] employs layer-wise quantization to minimize the increase in global loss. K-quant [15] is a group-quantization method that divides the weight matrix into multiple blocks, performing min-max quantization within each block. Due to outliers and dequantization overhead, most of the quantization algorithms focus on weight quantization only. However, by treating outliers carefully and implementing effective kernels, SmoothQuant [43] and OdesseyLLM [26] have tried weight-activation co-quantization to further lower the costs of LLM inference.
To optimize hardware resource utilization and accelerate inference, operator fusion and key-value cache are commonly used for on-device LLMs. Operator fusion [26] significantly improves computational efficiency by reducing the number of operations and eliminating the need to copy intermediate results between operations. Besides, as the inference process of generative LLM is autoregressive, implementing kv-cache can significantly reduce the computational cost by reusing key and value vectors across different positions. Although this increases memory requirements, the amount of cache space used is manageable due to single-batch inference and the relatively short context length typical of mobile applications. In addition, to address the quadratic cost of attention, MQA [37] and flash-attention [11] have optimized existing algorithms to accelerate inference by minimizing the computational overhead involved in attention calculations.
The researchers have attempted to develop lightweight LLM models with less than billion parameters and optimized them with these resource-efficient algorithms, which facilitate local LLM deployment. For example, Meta’s Llama 2 7B and Llama 3 8B models [12], after instruction fine-tuning, show effectiveness across various tasks. Built upon a similar block structure as Llama-2 but more lightweight, phi-3 models [33] introduced by Microsoft including phi-3-mini with 3.8B parameter and phi-3-small with 7B parameter achieves a quality that seems on-par with GPT-3.5. Moreover, Google’s Gemini Nano [16], specifically designed for mobile devices, comes in 1.8B and 3.25B sizes, which have been integrated into Pixel 8 Pro and can perform text summarization and provide intelligent responses locally.
2.2 On-device Deployment of LLMs
Traditional machine learning (ML) and deep learning (DL) models have established a rich and mature ecosystem for mobile platforms. For example, TFLite, a DL inference engine designed for mobile devices, offers a wide range of highly optimized operators. Additionally, some vendors such as Qualcomm have developed hardware-specific DL libraries, which enhance the efficiency of model deployment and inference. However, unlike convolutional models, which typically have only a few hundred MB of parameters, LLMs feature orders of magnitude more parameters, more complex architectures, and higher computational demands. This complexity necessitates the development of specialized engines or operators specifically designed for efficient LLM inference.
There are some inference frameworks widely used including llama.cpp [14], MLC LLM [38], mnn-llm [3], mllm [44], etc. Among them, llama.cpp and mnn-llm are optimized for LLM inference on CPUs, while MLC LLM leverages the powerful computational capabilities of GPUs. Despite the availability of dedicated AI accelerators such as NPUs, APUs, and TPUs in the latest generation of mobile SoCs [20], developing for these accelerators presents challenges due to the typically closed-source nature of vendor-specific SDKs. Currently, only mllm and PowerInfer-2 [46] claim to support Qualcomm NPUs for accelerating LLM inference on mobile devices. The development of LLM solutions for AI accelerators with dedicated architectures is still an ongoing challenge for developers.
In addition to those open-source LLM inference engines, mobile SoC vendors and manufacturers have also realized the importance of on-device LLM inference. Qualcomm, the vendor of Snapdragon SoCs, asserts that it can accelerate Llama-2 7B and Llama-3 8B models using the NPU on Snapdragon Gen2 and Gen3 platforms [36]. Similarly, MediaTek has announced optimizations for Llama 2 Generative AI applications using its APU on the Dimensity 9300 and 8300 [31]. However, both solutions remain closed-source and have yet to be validated through real-world applications.
2.3 Existing Measurement Studies
Several studies have focused on the measurement and evaluation of mobile-optimized LLMs. [25] and [27] both develop comprehensive benchmark suites to gain insights into various performance metrics, including latency, memory footprint, and throughput, as well as the impact of model size and quantization precision on accuracy. [25] considers a range of edge devices, from smartphones to Nvidia Jetson. In addition to throughput, it also investigates device power consumption over time. [27] primarily compares the performance of 22 lightweight models, offering a detailed analysis on model structure and operators.
However, none of these studies provide in-depth analysis about the impact of mobile hardware during inference, such as accelerator utilization, memory bandwidth, and fluctuations in CPU frequency. Although [27] touches on the impact of SoCs, it only examines the macro differences of SoCs from a single vendor. In contrast, our approach involves testing several SoCs from multiple vendors to gain a comprehensive understanding of how mainstream mobile processors support LLMs. This will help us identify hardware bottlenecks and potential future upgrades necessary for optimizing LLM performance on mobile devices.
Device | SoC | CPU | GPU | Avail./Total RAM | Type | OS | Release |
XiaoMi14 Pro | Snapdragon 8 Gen3 | Cortex-X4/A720/720/A520 | Adreno 750 | 12GB/16GB | top-tier | HyperOS | 2023.10 |
XiaoMi Pad6 Pro | Snapdragon 8+ Gen1 | Cortex-X2/710/510 | Adreno 730 | 12GB/16GB | top-tier | Android13 | 2022.5 |
Huawei Matepad11 Pro | Snapdragon 870 | Cortex-A77/A77/A55 | Adreno 650 | 5.5GB/8GB | mid-tier | Harmony 4 | 2021.1 |
Vivo Pad3 Pro | Dimensity 9300 | Cortex-X4/X4/A720 | Mali-G720 Immortalis MP12 | 10GB/16GB | top-tier | Android 14 | 2023.11 |
Huawei Matepad12.6 Pro | Kirin 9000E | Cortex-A77/A77/A55 | Mali-G78 MP22 | 8.5GB/12GB | high-tier | Harmony 4 | 2020.10 |
Huawei Nova7 | Kirin 985 | Cortex-A76/A76/A55 | Mali-G77 MP8 | 4GB/8GB | mid-tier | Harmony 4 | 2020.4 |
3 Experimental Setup and Methodology
3.1 Experimental Setup
Testing Devices. Table 1 shows the devices used in our measurement. There are mobile devices from three SoC vendors: Snapdragon, MediaTek, and Hisillicon. All CPUs feature eight Arm Cortex cores but differ in implementation details such as big.LITTLE configurations and cache sizes. It is noteworthy that the Snapdragon 8 Gen3 and Dimensity 9300 represent the most advanced SoCs available for mobile devices. We also test devices ranging from high-tier to mid-tier to capture a broad spectrum of hardware heterogeneity. In terms of GPU, the vast majority of mobile GPUs are from Adreno (Qualcomm) and Mali (Arm). Thus, we designate Snapdragon SoCs as representatives of Adreno GPUs and the rest as representatives of Mali GPUs.
Testing Model. We select Llama-2 B as a representative model for mobile-compatible LLMs. Although most mobile devices have RAM capacities ranging from GB to GB, the minimum effective available memory is around GB due to the substantial portion occupied by the operating system. Consequently, we determine that the B model size is the upper limit that most mobile hardware can support. Additionally, we employ the -bit quantization scheme, which is widely used and provides an optimal balance between model size and accuracy.
Frameworks and Engines. In this paper, we would like to reveal the performance of LLM on a variety of different accelerators. To achieve this, we select two representative frameworks: llama.cpp for CPU deployments and MLC LLM for GPU deployments, to assess the inference capabilities of general accelerators. For novel AI accelerators (e.g., NPU), rather than conducting our own device deployments, we refer to existing studies for results of mllm and PowerInfer-, which are only available for Qualcomm NPUs. These references will be discussed in subsequent sections.
llama.cpp [14] is a pure C language project based on the GGUF machine learning tensor library. The model weights are quantized into lower accuracy by K-quant method and stored as binary files in GGUF format. After compiling the code into a specific back-end executable file, users can try different models for inference and only need to download the corresponding model weight. There is no need to rely on other environments.
MLC LLM [38] is a high-performance LLM deployment engine built on the TVM machine learning compiler. Unlike llama.cpp, which focuses on optimizing for mobile CPUs, MLC LLM is designed to harness the power of mobile GPUs. Leveraging TVM, MLC LLM can automatically generate and optimize operators for a specific model and backend. These operators are then loaded and executed via the TVM Runtime. For Android-based GPUs, MLC LLM utilizes the OpenCL backend to ensure efficient execution.
Prompts and Outputs. To evaluate LLM performance in real-world scenarios, we use two types of prompts: a short prompt with tokens and a long prompt with tokens. The short prompt simulates typical user queries, such as asking for the meaning of a phrase, while the long prompt is designed for scenarios requiring more context, such as summarizing a news article. For generation, llama.cpp has fixed output lengths of tokens for -token prompts and tokens for -token prompts. In contrast, MLC LLM supports variable output lengths, allowing for generation up to tokens.
Measurement toolkits. In order to collect fine-grained hardware metrics such as real-time CPU frequency and memory bandwidth utilization, we choose Perfetto [1], Snapdragon Profiler [35] and Arm streamline [5] as our primary tools for monitoring hardware dynamics during inference. Perfetto offers comprehensive system-wide performance traces from Android devices, including data from the kernel scheduler, userspace instrumentation, and other sources. For CPU monitoring, Perfetto captures frequency, utilization, and event scheduling for each CPU core. Snapdragon Profiler and Arm Streamline are vendor-specific profilers used to track GPU behavior, including arithmetic and load/store unit utilization. Specifically, Snapdragon Profiler provides detailed timelines of OpenCL kernel executions, which is valuable for operator-level analysis.
Evaluation Metrics. We collect results from two main aspects. First, we gather performance metrics that are of interest to users, including prefill and decoding speed, memory footprints, and battery consumption. These results are present in Section 4. Second, we analyze hardware dynamics, focusing on real-time processor utilization and working frequency, impact of Dynamic Voltage and Frequency Scaling (DVFS) and operator implementation, which developers are concerned about. These results are present in Section 5. They provide clues for the performance metrics and can help us identify potential hardware and operating system bottlenecks during inference.
3.2 Workflow and Methodology
Workflow. Figure 1 shows the overall workflow how we deploy and evaluate models on mobile devices. The entire workflow is divided into four steps:
Step 1: Compile the model into a specific model library or binary executable file that comprises the inference logic.
Step 2: Transfer the library or executable file, model weights, and other necessary files to the test device and complete the deployment.
Step 4: Collect results such as inference latency, CPU frequency fluctuations, and other key metrics.
For llama.cpp, we build the model library and compile the source code into an executable file on each device individually, using CMake and the Android NDK. All the compile options are kept the same by default. When necessary, we adjust the compile options to enable specialized instructions to ensure optimal performance (Section 5.1.2). Since it relies on basic C libraries, no additional environment setup is required. After transferring model weights in GGUF format to the device, inference can be performed easily using the ADB command-line tool. For MLC LLM, there is a native application which TVM runtime and necessary libraries are packed in. The APK can be installed on the device, allowing interaction with the LLM through a graphical interface. MLC LLM cross-compiles the LLM models for the mobile platform, and on all devices, the runtime version including tvm and java is the same.
Measurement Methodology. To ensure accurate results, we perform two rounds of tests. In each round, a test is repeated five times, with an interval of seconds between each test. After the first round, the device reboots and rests for minutes to avoid overheating. This approach minimizes interference from other processes and reduces the impact of DVFS. The results are then averaged to provide a reliable measure of performance.
4 Performance: Users’ Perspectives
In this section, we present performance metrics that affect user experience, including token throughput, memory footprint, and energy consumption. We specifically compare the key performance metrics across different types of processors, including CPUs, GPUs, and specialized AI accelerators (NPUs).
4.1 Token Throughput and Memory Footprints
4.1.1 Performance on CPUs
Figure 3 shows the performance of CPUs, including prefill speed and decoding speed with llama.cpp. Overall, there has been a consistent improvement in both prefill and decoding speeds with new-generation processors. The Dimensity on Vivo Pad3 Pro, being the most advanced new-generation SoC, demonstrates the highest performance, achieving over speedup in prefill and nearly speedup in decoding compared to the older SoC Snapdrgon on the Huawei Matepad Pro. The Snapdragon 8 Gen (XiaoMi Pro) ranks second among all tested devices, showing % throughput of the Dimensity . Comparing Snapdragon SoCs across different tiers, each successive generation shows approximately a improvement in prefill speed, while decode speed improves even more significantly, ranging from to . We note that Snapdragon 870 (on Huawei Matepad11 Pro) and Kirin 9000E (on Huawei Matepad12.6 Pro) consist of the same cores, and the performance gap may result from higher CPU frequency on Kirin 9000E. Notably, a significant performance gap persists between Kirin SoCs and those from Snapdragon and MediaTek in terms of supporting LLMs.
Our results highlight a notable performance disparity in LLM inference across mobile CPUs. Specifically, top-tier SoCs like the Dimensity demonstrate significantly better LLM support compared to mid-tier SoCs. This suggests that for optimal LLM performance, developers should focus on leveraging the most advanced processors available. Furthermore, top-tier SoCs support advanced machine instructions that can significantly boost performance. We will examine this potential in detail in Section 5.1.2.
We further evaluate the performance with two scenarios: one involving a prompt of tokens with a generation of tokens, and the other involving a prompt of tokens with a generation of tokens. The larger input matrix in the latter case implies greater computational intensity and resource demands. Figure 3 illustrates that inference speed decreases as the token length increases. The results may be attributed to the impact of DVFS. With longer prompts and generation tasks, the increased load causes the CPU frequency to throttle more aggressively. Across all tested devices, the performance degradation for both prefill and decoding is relatively consistent, ranging between and .
In addition to latency, memory footprint is a also crucial metric. On all six devices, memory usage remains consistently around GB and shows minimal variation during inference once the model weights are loaded into main memory from external storage (e.g., Universal Flash Storage, UFS). This consistency is due to the fact that llama.cpp applies for a fixed-size memory block during initialization, which remains constant once allocation is complete. The allocated memory is then segmented into various components for different variables, including model weights, KV-cache, buffers, context and etc.
4.1.2 Performance on GPUs
Unlike CPUs, which are all from Arm, GPUs across vendors may have different architectures. Thus, it is challenging for developers to optimize for different GPUs. Figure 3 illustrates the end-to-end latency of MLC LLM across four high-tier GPUs from Arm Mali and Qualcomm Snapdragon. Note that Huawei Matepad11 Pro and Huawei Nova7 are excluded from this comparison due to subpar performance on devices with low RAM (GB). Among the tested GPUs, Mali GPUs, particularly the Mali-G78, exhibit significantly slow prefill speeds. Consequently, we only consider the performance with 64-token prompts. For the Mali-G78, due to its extremely poor prefill performance, we only present results for 32-token prompts as a reference.
The Adreno GPU consistently outperforms the Mali GPU in overall performance. Notably, Mali GPUs show unusually poor prefill performance. As for decoding, the Adreno delivers faster decoding speed than the Mali-G. Similarly, the Adreno achieves faster decoding speed compared to the Mali-G. Although Mali GPUs have poor prefill performance, Mali-G still achieves speedup in decoding than older-generation Adreno . Memory usage across the GPUs is nearly identical at around GB, with a slight increase to GB observed on the Vivo Pad3 Pro.
4.1.3 Performance on Specialized AI Accelerators
Recent years have seen the introduction of specialized AI accelerators in mobile SoCs, such as the Snapdragon Hexagon NPU, Kirin NPU, and MediaTek APU. However, these accelerators face several challenges, including limited support for dynamic shapes and floating-point operations, which complicates their use for LLM inference [46]. The Hexagon NPU, being the only mobile NPU with an open instruction set architecture, has garnered significant research interest. Successful deployments of mobile LLM models on the Snapdragon 8 Gen 3 have been reported. We summarize the findings from mllm, PowerInfer-, and Qualcomm AI Hub in Figure 4.
Qualcomm AI Hub, the official tool developed by the manufacturer, demonstrates a remarkable prefill speed of tokens per second, representing a -fold increase compared to both CPU and GPU. Similarly, mllm and PowerInfer-2 achieve substantial improvements in prefill speed, with an order of magnitude improvement in prefill speed. However, the decoding speed remains just slightly better than that of CPU and GPU based solutions. As we all know, prefill is compute-bound and decoding is memory-bound. This distinction likely accounts for the significant improvement in prefill speed observed with the NPU, whereas its decoding speed remains comparable to that of CPU/GPU.
4.1.4 Comparisons and Summary
1.
Given that mobile CPUs are built on the same Arm architecture and share common instruction sets, deploying LLMs on CPU is more straightforward and does not require specialized customization. As a result, CPUs have emerged as a preferred option for LLM inference on mobile platforms. However, our results show that they are limited in performance and thus not the optimal solution.
2.
GPUs, despite their theoretical advantages for LLMs due to their parallel float computing capabilities, do not show a clear advantage in practice. What’s counterintuitive is that MLC LLM on GPUs performs worse in prefill speed compared to llama.cpp on CPUs, a phenomenon that we will provide a detailed analysis in Section 5.1.3.
3.
In contrast, early trials with the Hexagon NPU demonstrate a significant advantage for specialized AI accelerators in inference tasks. However, due to their limited support for operators, adapting models and algorithms to specialized AI accelerators remains a challenging task for developers.
4.2 Battery Consumption
While the end-to-end latency of LLM inference on mobile CPUs shows promising potential, power consumption is also a key factor in determining the feasibility of these models for local, always-on user access.
Table 2 compares the power consumption and inference speed between the Vivo Pad3 Pro and Huawei MatePad 12.6 Pro. To obtain the power consumption, we read the data provided by the OS settings, which records and displays the power consumption of a certain application. Old-generation devices are excluded from this comparison due to significant differences in battery capacity and power-saving policies compared to new-generation hardware. We perform consecutive rounds of inference on the CPU, with a prefill length of and a generation length of . The Dimensity 9300 exhibites a speedup in prefill and a speedup in decoding, while consuming only of the power used by the Kirin E. These results suggest that top-tier SoCs are capable of delivering substantial energy savings alongside improved performance.
Most modern smartphones feature battery capacities between mAh and mAh. At a power consumption rate of mAh per inference round, a device could theoretically support over rounds of inference. This aligns with the findings in [25], reinforcing the feasibility of LLM inference on mobile devices. However, a significant rise in device temperature during inference suggests that LLM inference is a power-intensive task. We recorded an increase from °C to °C during a single inference on Huawei Matepad12.6 Pro. Thus, optimizing the power efficiency of LLMs on mobile devices is crucial for sustainable and long-term deployment.
Device | Power Drain | Prefill | Decode |
(mAh/round) | (token/s) | (token/s) | |
Vivo Pad3 Pro | |||
Huawei Matepad12.6 Pro |
5 Performance: Developers’ Perspectives
In this section, we present results that developers care about, focusing on CPU/GPU utilization, DVFS and scheduling strategies. We also investigate the impact of different inference frameworks. We hope that these results can help developers identify bottlenecks in LLM inference and ultimately lead to improvements in system performance.
5.1 Hardware Capabilities
While upgrading hardware can enhance LLM local inference performance, it is also crucial to assess whether we are fully utilizing the capabilities of existing hardware. To address this, we use specialized profilers to monitor and capture dynamic utilization of the CPU and GPU during inference. This allows us to explore the potential of current hardware and identify opportunities for further accelerating LLM inference.
5.1.1 Multi-threads on CPU Cores
SoC | Snapdragon Gen | Dimensity | Snapdragon + Gen | Kirin E |
Prime | Cortex-X4 (GHz) | Cortex-X4 (GHz) | Cortex-X2 (GHz) | Cortex-A77 (GHz) |
Performance | Cortex-A720 (GHz) Cortex-A720 (GHz) | Cortex-A720 (GHz) Cortex-A720 (GHz) | Cortex-A710 (GHz) | Cortex-A77 (GHz) |
Efficiency | Cortex-A520 (GHz) | - | Cortex-A510 (GHz) | Cortex-A55 (GHz) |
Most popular mobile SoCs utilize the "big.LITTLE" architecture for their CPUs, which balances performance with power efficiency. This configuration typically includes multiple cores organized into two distinct clusters: "big" (prime and performance cores) and "little" (efficiency cores), as illustrated in Table 3. While it is commonly assumed that high-load tasks are best handled by the "big" cores for optimal performance, our tests reveal that the ideal core configuration can vary across the two stages of LLM inference.
We evaluate three core configurations: using only big cores, a combination of big and little cores, and all available cores. To enable parallel inference across multiple CPU cores, we adjust the number of running threads. Since inference threads are prioritized to run on big cores, the number of running threads implies which cores are active. For instance, on the Snapdragon 8 Gen 3 (which has six big cores), six threads correspond to all big cores, while on the Snapdragon 8+ Gen 1 (which has only four big cores), the same configuration includes all big cores and two little cores.
Figure 5 illustrates how these core configurations impact inference speed across the four tested devices. During the prefill stage, prompt processing speed is primarily driven by the big cores. The contribution of the little cores to acceleration depends on the performance gap between the big and little cores. On devices with limited computational power, additional efficiency cores can help accelerate inference. For instance, utilizing all cores delivers the best performance on the Kirin 9000E. However, for devices with more powerful big cores, adding little cores can actually reduce performance. For example, on the Snapdragon 8 Gen 3, incorporating two little cores results in a slowdown in inference speed. This highlights the importance of optimizing core configurations based on each device’s specific CPU capabilities to maximize efficiency. Interestingly, the Dimensity 9300 benefits from using all cores because of its All-Big-Core architecture [30].
In the decoding phase, adding more little cores typically leads to a decline in performance. For instance, both the Kirin 9000E and Snapdragon 8+ Gen 1 achieve their peak decoding speeds with four active big cores. This indicates that the maximum decoding speed on a device is largely dictated by the performance of the big cores. Unlike the prefill stage, decoding is more heavily constrained by memory bandwidth. This is evident in the fact that adding more big cores results in a smaller performance improvement compared to prefill, while incorporating little cores leads to a more noticeable performance drop.
5.1.2 Speedup with Special Machine Instructions
CPUs are generally more efficient at handling integer (INT) operations compared to floating-point (Float) computations. This performance advantage in INT operations stems from two key factors: first, CPUs typically offer higher instruction throughput for INT operations than for floating-point ones. Second, for matrix multiplication tasks, many CPUs support specialized INT8 instructions, which further accelerate these computations.
To further explore the potential of Arm CPUs, we test the Llama-2-7B model on top-tier SoCs and recompiled llama.cpp directly on the device with the i8mm flag to enable INT8 matrix multiplication instructions. This ensures that the machine code consists of smmla[9] and sdot[7]. The sdot instruction accelerates vector computations by processing multiple INT8 dot products simultaneously, making it well-suited for matrix-vector multiplication during the decoding. In contrast, the smmla instruction handles block matrix multiplication and delivers twice the multiply-accumulate efficiency of sdot, making it ideal for accelerating matrix operations in the prefill stage.
We note that Arm developers have proposed the prearrangement of weights in blocks to avoid pseudo-scalar operations and fully exploit the parallel capabilities of specialized instructions [18]. In the original weight layout, only a single column of weights is processed at a time, and it requires multiple dot product operations. The optimized layout, however, distributes the weight columns across multiple computational lanes, enabling parallel processing of several weight columns in a single instruction. This approach not only increases throughput but also reduces memory access by sharing input data within lanes, further enhancing inference efficiency.
Figure 6 highlights a notable performance boost. When focusing on peak inference performance, prefill speed is accelerated by under the same configuration. This speedup is primarily achieved by increasing throughput, allowing more computations to be processed within a single instruction. On the Snapdragon 8 Gen 3 (Xiaomi 14 Pro), adding all efficiency cores (little cores) results in a decline compared to using only the big cores. This suggests that the lower frequency and IPC (instructions per cycle) of efficiency cores may hinder performance during high-demand tasks.
GPU Name | FP16 | Est./Theoretical Max |
(GFLOPS) | Memory Bandwidth (GB/s) | |
Mali-G78 MP22 | /- | |
Mali-G720 Immortalis MP12 | /77 | |
Adreno 730 | /- | |
Adreno 750 | /77 |
5.1.3 Utilization of GPU units
We note an intriguing exception regarding performance on Mali GPUs. The Mali-G Immortalis MP12, found in the Dimensity 9300, has better hardware parameters but has poorer performance than Adreno in Snapdragon SoCs. We measure the maximum throughput for FP16 operations on each device using clpeak[24]. It is a benchmarking tool intended for calculating GPU performance by executing different OpenCL kernels on a particular device. We also measure the actual maximum memory bandwidth using clpeak and present the maximum theoretical bandwidth in Table 4. From Table 4, we can see that the Mali-G720 Immortalis MP12 exhibits times the throughput of the Adreno in float16 operations. However, in Figure 3, it shows a prefill speed that is times slower.
The possible reason for this discrepancy comes from the GPU utilization. Figure 7 presents the average utilization and memory bandwidth during inference on Mali-G720 and Adreno 750. It shows that MLC LLM does not efficiently utilize the Mali GPU ALUs. During the compute-bound prefill stage, the Mali-G720’s arithmetic unit utilization averages less than , while the Adreno achieves around . Similarly, in the memory-bound decoding stage, the Mali-G achieves a real-time memory read bandwidth of GB/s, which is only roughly two-thirds of the GB/s of the Adreno 750. These results indicate existing LLM implementations cannot fully utilize the available computing resources on mobile GPUs. It implies that while OpenCL serves as a general-purpose backend, it is essential to implement GPU-specific kernels to fully leverage the hardware capabilities.
5.2 Impact of DVFS
In Section 4.1.1, we observed performance degradation as the length of the prompt and generation increased. This issue relates to the system’s ability to handle long contexts and maintain consistent inference performance across multiple rounds of dialogue. Although previous studies, such as [25] and [27], have discussed potential causes like Dynamic Voltage and Frequency Scaling (DVFS) and thermal throttling, they lack detailed analysis on how DVFS specifically impacts inference performance and whether its effects differ across hardware and operating systems. To address this, we analyze real-time variations in CPU frequency and utilization during 20 continuous inference tests on four devices with warm up, covering different operating systems (Android, HyperOS, HarmonyOS) and vendors (MediaTek, Qualcomm, Hisilicon).
Figure 9 illustrates a general decline in performance as the rounds of inference increase. Specifically, on the Snapdragon Gen (on Xiaomi14 Pro), latency increases by up to , dropping from tokens/s to tokens/s. In contrast, Dimensity and Kirin E exhibit only minor latency fluctuations, with a maximum decrease of approximately . This trend aligns with the data presented in Figure 3, which shows that the Xiaomi Pro experiences a more significant performance drop as prompt length increases.
Figure 9 shows the frequency variation on the prime CPU core at different rounds of inference. It reveals that Snapdragon SoCs (e.g.Huawei Matepad11 Pro, Xiaomi14 Pro) exhibit a more aggressive DVFS policy. During subsequent stages of the inference, both prime and performance cores on Snapdragon 8 Gen 3 experience a rapid reduction in frequency. By the th inference round, the frequency is nearly halved compared to its initial value. To investigate the factors influencing DVFS policy, we further compare two Huawei tablets equipped with identical CPU cores but different SoCs—one powered by Snapdragon (Snapdragon 870 on Huawei Matepad11 Pro) and the other by Kirin (Kirin 9000E on Huawei Matepad12.6 Pro). The results indicate that the vendor plays a significant role in shaping the DVFS policy. For instance, although the maximum frequencies of both devices are GHz, the Snapdragon 870 typically keeps the prime core frequency below GHz, while the Kirin 9000E sustains a higher frequency of around GHz during inference.
5.3 Model Preparation Latency
In Section 5.2, we have studied the inference performance with warm-up. However, in real-world applications, users may trigger the model sporadically, leading to scenarios where the model must perform a cold start without a warm-up, known as the cold start. Thus, the time required to generate the first token becomes critical to user experience. The process of generating the first token involves several stages: loading model weights from external storage into memory, preparing memory for variables, and executing prefill and decoding. In this section, we focus on the preparation procedure before prefill and evaluate the latency across various devices. We perform tests on six devices and measure the time spent on preparation with and without warm-up. If the device is tested without warm-up, we reboot it to refresh RAM and cache. Results are present in Figure 10.
For the llama.cpp running on CPUs, we compare the time before prefill of the first token in both cold-start and warm-up scenarios. The Dimensity 9300 (on Vivo Pad3 Pro) delivers the best overall performance, with an average model loading time of approximately ms. The Snapdragon 8 Gen 3 (on Xiaomi14 Pro) ranks second, with its worst-case performance comparable to that of the Dimensity 9300. We recognize that the warm-up process significantly reduces preparation time for llama.cpp, primarily due to the use of mmap, which allows the inference process to directly access model weight files from disk. The mmap function accelerates loading by eliminating the need to copy data between user and kernel space. Additionally, the mapping remains valid for other processes accessing the same file, meaning that after a warm-up, the model weights are already loaded into memory. This allows subsequent inference processes to use the preloaded weights, thereby reducing latency. However, this benefit depends on the memory retaining the loaded weights—if they are overwritten by other files, the warm-up advantage is lost. From Figure 10, we observe that all Huawei devices exhibit unusually long preparation times during cold starts. On Huawei devices with Kirin SoCs, the delays are even more pronounced, with the worst loading times exceeding seconds. Huawei Matepad 11 Pro, despite using a Snapdragon SoC, also experiences extended delays. We identify the underlying causes for these exceptions are related to the mmap function in Harmony OS and we will discuss it in Section 5.4.1.
In contrast, for MLC LLM running on GPUs, no significant improvement is observed after warm-up, with loading times remaining approximately five times longer on Adreno GPUs and seven times longer on Mali GPUs compared to llama.cpp (with warm-up). This is due to mobile system architecture. Although the CPU and GPU on mobile devices share the same physical DRAM, they operate in separate memory spaces and cannot directly access each other’s data[47]. As a result, model weights must be copied from the CPU memory space to the GPU memory space, introducing additional latency. Specifically, MLC LLM uses OpenCL kernels to transfer data between the CPU and GPU via buffers, which contributes to the delay.
Type | Memory(RAM) | Storage(ROM) | L3 Cache |
XiaoMi14 Pro | 4800 MHz(LPDDR5X) | UFS 4.0 | 12MB |
Vivo Pad3 Pro | 4800 MHz(LPDDR5X) | UFS 4.0 | 18MB |
The variations in preparation times across devices are primarily linked to memory bandwidth and UFS speed. For example, the preparation times for Xiaomi14 Pro (Snapdragon Gen ) and Vivo Pad3 Pro (Dimensity ) are comparable on both CPU and GPU. However, the subtle hardware differences still influence performance. Snapdragon 8 Gen 3 shows a slight advantage on the GPU, while the Dimensity performs better on the CPU. As shown in Table 4 and Table 5, the Snapdragon Gen (Adreno 750) benefits from higher real-time GPU memory bandwidth, whereas the Dimensity 9300’s larger L3 cache allows for more efficient reuse of loaded weights after the warm-up. verall, preparation times on both CPUs and GPUs across different vendors are short enough to meet the demands of most real-time applications.
5.4 Inference Engines
In the previous sections, we studies the impact of hardware on LLM inference. However, some performance discrepancies stem from the design and implementation of inference engines rather than inherent hardware capability limitations. In this part, we will focus on the role of inference engines and provide recommendations for framework developers to improve performance.
5.4.1 llama.cpp
The first issue concerns the exceptionally long model preparation time without warm-up on Huawei devices. Through timestamp analysis, we identify that the primary cause is the execution of the mmap function. On the Huawei Matepad 12.6 Pro, mmap execution takes up to seconds, whereas after warm-up, this time is drastically reduced to seconds. This suggests that the delay arises from the memory mapping process rather than loading from UFS. We further verify this by testing without file mapping, which reduces the average preparation time to seconds. Additionally, on the Huawei Matepad 11 Pro with a Snapdragon SoC, the average preparation time is seconds—showing noticeable improvement over other Huawei devices but still significantly slower than the non-Huawei devices. These findings suggest that the issue may be related to mmap behavior on HarmonyOS, possibly exacerbated by Kirin SoCs.
Secondly, we compare different versions of llama.cpp on the Snapdragon 8 Gen 3, Dimensity 9300, and Kirin 9000E in Figure 11 to evaluate how code optimizations affect performance across various SoCs. The results show that new-generation SoCs such as Snapdragon 8 Gen 3 and Dimensity 9300 experience significantly greater performance gains compared to others. Additionally, the latest version of llama.cpp optimizes model weight arrangement to align with INT8 matrix multiplication instructions, resulting in a peak prefill speed of up to 56 tokens/s. This improvement highlights that to better harness the full capabilities of processors, developing support for these advanced instructions is necessary.
5.4.2 MLC LLM
In Section 5.1.3, we explored the factors contributing to the suboptimal performance of MLC LLM on non-Snapdragon SoCs. While we attribute this issue to insufficient hardware unit utilization, it is important for developers to understand what implementation optimizations can address this inefficiency. We observe that MLC LLM supports local execution on devices like the Orange Pi [39], which features a Mali-G610 GPU. Despite the fact that Mali-G’s theoretical computation capability being significantly lower than the Mali-G on the Dimensity 9300, the prefill speed on the Orange Pi is tokens per second—more than twice that of the Mali-G. This discrepancy suggests that theoretical computation capability alone is not sufficient to explain performance differences and points to the importance of optimization. Furthermore, MLC LLM configurations vary widely across target devices. For instance, the loop unroll count is set to on Adreno but for Mali, underscoring the necessity of optimizing operators for specific hardware accelerators. Given the greater heterogeneity of mobile GPUs compared to CPUs, achieving optimal performance requires tailored optimizations for each platform.
6 Implications
In this section, we investigate several aspects that can improve mobile LLM performance.
Memory Bottleneck According to Figure 6, the prefill speeding is times faster than decode. Decoding is primarily constrained by memory bandwidth. Current optimizations, such as adding more cores and using specialized instruction sets, have proven effective in accelerating prefill. However, these approaches offer limited improvements for decoding speed. To address this challenge, more effective solutions are needed. Developers should consider designing algorithms that optimize memory utilization during decoding.
Hardware Instructions A key aspect of SoC evolution is the continuous enhancement of instruction sets. New machine instructions often increase operand width, resulting in higher throughput. To fully harness the potential of hardware, developers must consider the characteristics of model operators and take advantage of hardware-accelerated instructions. This may also involve adjusting computational precision, such as quantizing activations to -bit to utilize instructions like smmla. In Section 5.1.2, we discuss the improvement brought by Armv8 instructions. For next-generation CPUs based on the Armv9 architecture, advanced vector instructions like SVE2[10]- which enables longer computation vectors-are supported. Furthermore, Arm’s SME instructions[8], specifically designed for matrix multiplication, present additional optimization opportunities. Staying current with these advancements is crucial for developers seeking to maximize inference speed on mobile platforms.
Faster First Generation The time required for generating the first token includes model preparation, prefilling and one round of decoding. Since the majority of model preparation involves I/O operations (loading model weights into memory), storing some of the weights in RAM can help reduce preparation time. As for prefill and decode, the CPU can run at a higher frequency when a new inference process starts, ensuring the fastest response.
7 Discussion
In this section, we explain the rationale behind our choices in this measurement study.
Large Lanuage Models. There are many lightweight LLMs, with parameters ranging from 1B to 13B, that can fit in memory on most high-end mobile devices. Additionally, strategies have been developed to run larger models by offloading part of the model weights to external storage. In this paper, rather than focusing on performance differences across models, we concentrate on the hardware capabilities to identify bottlenecks that limit LLM performance and thoroughly understand how much untapped potential remains.
Most existing generative AI models are transformer-based, where the primary operations are matrix-matrix or matrix-vector multiplications. While different models may vary in terms of matrix size or the number of transformer blocks, the core operations remain highly resource-demanding matrix computations. Understanding the hardware’s capacity to handle these operations is key to optimizing LLM performance on mobile devices.
Quantization and Model Accuracy. An important aspect of evaluating a model’s capability is the accuracy of the responses generated by the LLM. However, for on-device LLM inference, the factors influencing model accuracy are primarily subject to the model architecture and the quantization of its weights. Based on this fact, MobileAIBench[34] provides a comprehensive benchmark for analyzing accuracy across various NLP tasks under different quantization schemes for both LLMs and large multimodal models (LMMs). Since our primary focus is on hardware-related performance metrics, we have chosen the most widely used quantization scheme, which offers the best balance between speed and accuracy. While lower-bit quantization algorithms may enhance inference speed, we excluded them from this study due to their significant impact on accuracy.
Multimodality Models. Driven by the growing demand for advanced AI assistants, the deployment of large multimodal models (LMMs) on mobile devices is an inevitable trend. However, we excluded LMMs from our study for two main reasons. First, current LMMs may not yet be ready for mobile deployment. For example, MobileAIBench reports that testing Llava-Phi-2[48] on the iPhone 14 resulted in a first token generation of seconds. It remains unclear whether this poor performance is due to hardware constraints or inefficient operator implementation. Second, since LLMs and LMMs share significant similarities in both model architecture and operator implementations, the insights gained from on-device LLM tests can likely be extrapolated to large multimodal models.
8 Concluding Remarks
In this paper, we present a comprehensive measurement study of LLM inference on mobile platforms, offering insights into both user-experience metrics (such as token throughput and energy consumption) and critical hardware and system characteristics (including CPU/GPU utilization, DVFS, and file mapping latency). Our analysis reveals how hardware capabilities and system dynamics impact on-device LLM performance and highlights potential bottlenecks affecting mobile LLM applications. Additionally, we propose potential directions for enhancing on-device LLM inference performance. We hope that this study can provide insights for both the development of on-device LLMs and the design for future mobile system architecture.